Show all steps, thanks

Additional Problem 3. If X is a continuous random variable having cdf F, then its median is defined as that value of m for which F(m) 0.5. Find the median for random variables with the following density functions (a) f(x) = e-*, x 0 (b) f(x) = 1, 0 〈 x 1. (c) f(x) 6x(1 - x),0 <1.
Answer #1

a)

as F(m)=0.5

\int_{0}^{m} f(x) dx =0.5

\int_{0}^{m}e-x dx =(-e-x) |m0 =0.5

1-e-m =0.5

m=-ln(0.5)=0.6931

b)

\int_{0}^{m} f(x) dx =0.5

\int_{0}^{m}1 dx =(x) |m0 =0.5

m=0.5

c)

\int_{0}^{m} f(x) dx =0.5

\int_{0}^{m}6(x-x2) dx =(6(x2/2-x3/3) |m0 =0.5

-2m3+3m2-0.5 =0

solving above: m=0.5